OpenStudy (anonymous):

What is the radius of of a circle with a circumference of 32π yd? A. 4 yd B. 8 yd C. 16 yd D. 32 yd

OpenStudy (anonymous):

PLZ HELP!!! FAN AND MEDAL!!!!

OpenStudy (anonymous):

to best answer

OpenStudy (bradely):

circumference,2πr=32π 2r=32 r=16 yd Source: http://www.mathskey.com/question2answer/

OpenStudy (johnweldon1993):

\[\large C = 2\pi r\] \[\large 32 \cancel{\pi} = 2\cancel{\pi} r\] \[\large r = \frac{32}{2} = ?\]

OpenStudy (anonymous):

thank you @johnweldon1993 and @bradely

OpenStudy (johnweldon1993):

anytime!

OpenStudy (anonymous):

what grade are you in @johnweldon1993

OpenStudy (johnweldon1993):

I'm a junior in university @CUTIEGIRL01